OpenStudy (anonymous):

Find the least common denominator for these two rational expressions. s^3/s^2+2s+1 and 4/s^2+6s+5

OpenStudy (anonymous):

can someone help.

OpenStudy (anonymous):

Try factoring the denominators

OpenStudy (jdoe0001):

hmmm actaully one sec.. the 2nd one is off =)

OpenStudy (jdoe0001):

\(\large {\cfrac{s^3}{s^2+2s+1}\quad \cfrac{4}{s^2+6s+5} \\ \quad \\ hint:\quad \begin{array}{lllllll} s^2&+2s&+1&\quad &s^2&+6s&+5\\ &\uparrow&\uparrow &&&\uparrow&\uparrow\\ &1+1&1\cdot 1&&&5+1&5\cdot 1 \end{array} }\) factor them out, see what you get

OpenStudy (anonymous):

(s+1)^2 and (s+5)(s+1)

OpenStudy (jdoe0001):

yeap thus \(\bf \cfrac{s^3}{s^2+2s+1}\quad \cfrac{4}{s^2+6s+5}\implies \cfrac{s^3}{(s+1)(s+1)}\quad \cfrac{4}{(s+5)(s+1)} \\ \quad \\ LCD\implies (s+1)(s+1)(s+5)\) since the (s+1) is already included, from the 1st denominator, it can be skipped for the 2nd, thus

OpenStudy (jdoe0001):

but (s+1) has to be included twice, so it can divide the \((s+1)^2\) denominator